### v7

# Data-Copilot: Bridging Billions of Data and Humans with Autonomous Workflow

Abstract
Industries such as finance, meteorology, and energy generate vast amounts of heterogeneous data daily. Efficiently managing, processing, and displaying this data requires specialized expertise and is often tedious and repetitive. Leveraging large language models (LLMs) to develop an automated workflow presents a highly promising solution. However, LLMs are not adept at handling complex numerical computations and table manipulations and are also constrained by a limited context budget. Based on this, we propose Data-Copilot, a data analysis agent that autonomously performs querying, processing, and visualization of massive data tailored to diverse human requests. The advancements are twofold: First, it is a code-centric agent that receives human requests and generates code as an intermediary to handle massive data, which is quite flexible for large-scale data processing tasks. Second, Data-Copilot involves a data exploration phase in advance, which explores how to design more universal and error-free interfaces for real-time response. Specifically, it actively explores data sources, discovers numerous common requests, and abstracts them into many universal interfaces (code modules) for daily invocation. When deployed in real-time requests, Data-Copilot only needs to invoke these pre-designed interfaces, transforming raw data into visualized outputs (e.g., charts, tables, and text) that best match the user’s intent. Compared to generating code from scratch, invoking these pre-designed and compiler-validated interfaces can significantly reduce errors during real-time requests. Additionally, interface workflows are more efficient and offer greater interpretability than code. We open-sourced Data-Copilot with massive Chinese financial data, such as stocks, funds, and news, demonstrating promising application prospects. Quantitative evaluations indicate that, compared to direct code generation and other agent strategies, our exploration-deployment strategy addresses human requests more accurately and efficiently, i.e., with higher success rates and lower token consumption.
Figure 1: We present two strategies for using LLMs to automate data analysis. Upper: Directly using LLMs to read and process massive data is limited by context length, the LLM’s capabilities, and output format constraints. Bottom: Data-Copilot is a code-centric agent that utilizes code to handle extensive data analysis tasks. It actively explores how to design more universal and error-free interface modules, improving the success rate of real-time requests. Faced with real-time requests, Data-Copilot invokes self-design interfaces and constructs a workflow to fulfill human intent.
1 Introduction
In the real world, vast amounts of heterogeneous data are generated every day across various industries, including finance, meteorology, and energy, among others. Humans have an inherent and significant demand for data analysis because these wide and diverse data contain insights that can be applied to numerous applications, from predicting financial trends to monitoring energy consumption. However, these data-related tasks often require specialized knowledge and tedious manual labor.
Recently, the advancement of large language models (LLMs) (Zeng et al., 2023; Touvron et al., 2023; OpenAI, 2022), especially when combined with techniques such (Wei et al., 2022b; Kojima et al., 2022) has demonstrated their capability to handle complex tasks. Given the vast amounts of data generated daily, can we leverage LLMs to create an automated data processing workflow that performs data analysis and outputs visualization results that best match user expectations?
An intuitive solution is to directly use LLMs to read and process massive data (Wu et al., 2023c; Zha et al., 2023), just as we do with text. However, several challenges must be considered: (1) Due to the input context limitations of LLMs, it is challenging for LLMs to directly read and process massive data as they do with text. Besides, it also poses the potential risk of data leakage when LLMs directly access private data sources. (2) Data processing is very complex, involving many tedious numerical calculations and intricate table manipulations. LLMs are not adept at performing these tasks. (3) An LLM can only output text, whereas we usually prefer charts or tables when analyzing data. These challenges significantly constrain the application of LLMs in data-related tasks.
Recently, many works have explored alternative solutions. LiDA (Dibia, 2023) and GPT4-Analyst (Cheng et al., 2023) focus on automated data exploration. Sheet-Copilot (Li et al., 2023b), BIRD (Li et al., 2023c), DS-Agent (Guo et al., 2024) and DB-GPT (Xue et al., 2023) apply LLMs to data science domain like Text2SQL. Besides, numerous LLM-based agent methods have recently emerged (Wu et al., 2023a; Huang et al., 2023; Chen et al., 2023b; Hong et al., 2023; Wu et al., 2023b). Data Interpreter (Hong et al., 2024) proposes a code-based agent for data science through the plan-code-verify paradigm. These methods showcase the potential of LLMs in completing complex daily tasks through agent design paradigms.
Inspired by this, we advocate leveraging the coding capabilities of LLM to build a data analysis agent. Acting like a human data analyst, it receives human requests and generates code as an intermediary to process massive data and output multi-modal results (e.g., chart, table, text) for humans, reducing tedious human labor and eliminating the need for direct data access by the LLM. However, creating a reliable code agent for various data analysis tasks is also not an easy feat. Firstly, LLM-generated code often contains formatting errors, logical inconsistencies, and code hallucinations, such as fabricating non-existent functions, which may prevent it from being deployed in real-world environments. Secondly, most code-based agent generates code for solving user requests, without considering the relations between different requests and only addressing each issue independently.
Based on this observation, we propose Data-Copilot, an LLM-based agent with an innovative exploration phase to achieve more reliable data analysis. First, Data-Copilot is a code-centric agent that connects massive data sources and generates code to retrieve, process, and visualize data in a way that best matches the user’s intent. The code-centric design empowers it to efficiently and securely handle extremely large-scale data and nearly all types of data analysis tasks. Besides, Data-Copilot also incorporates a data exploration phase. During this phase, it actively explores how to design more universal and error-free interfaces (code modules) based on data schemas in advance. In a real-world deployment, Data-Copilot flexibly invokes pre-design interface modules to handle user requests, creating an interface workflow that transforms raw data into desired visualization results.
This design can bring three advantages. Firstly, this exploratory process allows Data-Copilot to analyze connections between data, design general interfaces for similar requests, and pre-validate their correctness, thereby reducing errors in real-time responses. Secondly, when faced with massive requests, our agent only needs to invoke these pre-designed interfaces rather than generate redundant code, significantly improving generation efficiency. Lastly, compared to lengthy code, these interfaces provide greater interpretability when solving user requests, since it is easier for human reading and interaction.
As shown in Figure 1, Data-Copilot contains three key steps for data exploration:
  1. 1.
    Explore data and synthesize requests: Data-Copilot discovers potential requests and a broader range of human need scenarios from data. It involves a "self-exploration" process to generate massive synthesized requests based on all data schemas and seed requests.
  2. 2.
    Interface design and test: Data-Copilot generates modular code for request solving. Each interface is a functional module, and test cases are generated to verify the designed interface.
  3. 3.
    Interface optimization: To improve the versatility of the interface, Data-Copilot merges similar interfaces and revises the erroneous interfaces based on the compiler’s feedback.
After the above exploration-design-optimization steps, Data-Copilot designs more general and error-free interfaces, e.g., data acquiring, forecasting, and visualizing modules, to accomplish data analysis tasks. When faced with real-time requests, Data-Copilot invokes these predefined interface modules to create a concise interface workflow for user requests. For different requests, Data-Copilot can flexibly deploy various invocation structures, such as step-by-step serial workflows, parallel, or loop. It can even output a hybrid form of interface workflows and raw code for some "unfamiliar" requests. Our contributions are threefold:
Figure 2: Overview of Data-Copilot. Data Exploration: First, it performs a self-exploration process to uncover potential human requests from data sources. Then it abstracts many universal and error-free interfaces from synthesized requests, including interface designing, testing, and optimizing similar interfaces one by one. This exploration-design-optimization process is operated in advance. Workflow Deployment: Upon receiving real-time requests, Data-Copilot can quickly invoke existing interfaces and deploy a workflow for familiar requests, or flexibly combine interfaces and raw code for "uncovered" requests.
  1. 1.
    We propose a code-centric agent, Data-Copilot, for massive data analysis tasks. It leverages LLM and its code generation capabilities to automate data querying, processing, and visualization tailored to human requests, reducing tedious human labor.
  2. 2.
    We decouple code generation into two phases: data exploration and workflow deployment. In the exploration phase, Data-Copilot learns to design universal, error-free interface modules tailored to data. These interfaces enhance the success rate and efficiency of real-time responses. In the face of real-time requests, it flexibly invokes these interfaces to address users’ diverse requests.
  3. 3.
    We open-source Data-Copilot for diverse Chinese financial data analysis, including stocks, funds, and live news. Quantitative evaluations indicate that our agent outperforms other agent strategies, with higher success rates and lower token consumption. Besides, interface workflows are more convenient for human inspection and interaction, offering stronger interpretability.
2 Data-Copilot
Data-Copilot is a code-centric agent that reads human requests, writes code, or invokes predefined interface modules to retrieve, process, and visualize data, returning results according to human intent. Data-Copilot operates in two phases: Data Exploration and Workflow Deployment. In the first phase (Section 2.1), Data-Copilot designs a self-exploration process to discover numerous potential needs based on data schemas (Step 1: exploration). Then based on these synthesized requests, it abstracts universal interface modules for each request (Step 2: design). After that, it optimizes similar interfaces and tests their correctness, ensuring each interface is correct and universal (Step 3: testing and optimization). The whole exploration-design-optimization process is operated in advance, yielding many generic, error-free interfaces for subsequent use.
When faced with real-time requests (Section 2.2), Data-Copilot can flexibly invoke pre-designed interfaces or directly generate raw code based for the user’s request. In most cases, existing interfaces can cover the majority of real-world requests, significantly enhancing both success rate and response speed. We provide a detailed prompt for two phases in Section F.1 and F.2.
2.1 Data Exploration
Let’s review how human data analysts operate. Initially, they need to observe what data is available, understand the data formats, and determine how to use them. Subsequently, humans often abstract generic modules to simplify the code logic and test these modules for usability. Similar to this, Data-Copilot also autonomously explores data and derives insights from vast data sources, including the relations between the data, and the potential requests associated with the data. Then Data-Copilot abstracts these exploratory insights into numerous reusable code components (interfaces), testing their correctness and optimizing their generality. This process of exploring data, identifying common requests, designing general interfaces, and testing and optimizing their performance is conducted in advance on its own.
Self-Exploration Request
To explore data and mine insights, we design a self-exploration process. Beginning with some seed requests collected from humans, LLMs are prompted to read the data and generate a large number of requests, each representing a potential demand scenario. This process is similar to (Wang et al., 2022a; Dibia, 2023), but the LLMs should generate requests specifically based on provided data. As shown in Figure 2, when the LLMs observe that the economic database contains historical GDP and CPI data, they generate multiple related requests, e.g., Compare the CPI of China and the USA..., or I want to see the GDP trend over the past ten years. Each request involves one or more types of data.
To achieve this, we first generate a parsing file for each data source to help LLM understand the data. Each file includes a description of the data, the access method, the data schema (the name of each column), and a usage example. Then we feed the parsing files and a few seed requests into the LLMs, and prompt LLMs to generate more diverse requests based on these data, covering different demand scenarios as comprehensively as possible. The brief prompt is shown as follows:
###Instruction: Given some data and its description, please mimic these seed requests and generate more requests. The requests you generate should be as diverse as possible, covering more data types and common needs.
###Seed Request: {request1, request2,...}
###Parsing file for GDP_Data:{
Description: This data records China’s annual and quarterly GDP...,
Access Method: pro.cn-gdp(start-time, end-time, frequency,...),
Output Schema: Return 9 columns, including quarter, gdp, gdp-yoy...,
Usage:{
Example: pro.cn-gdp(start-q=’2018Q1’, end-q=’2019Q3’,..),
First Row: {2019Q4, 990, ..}, Last Rows: {2018Q4, 900, ..}}}
###Parsing file for Stock_data:{...}, ....
How to ensure the quality of the generated requests? A common issue is that the LLM often proposes a request about the data that doesn’t exist. To address this, we design a backward verification process to check the rationality of these requests. Specifically, we instruct the LLM to reverse-convert the generated request into the desired data source and other key information, and then we verify the existence of such data, thereby filtering out hallucinatory requests. Upon manually checking some requests, we found that the generated requests generally met our expectations.
Interface Design
After generating massive requests through self-exploration, Data-Copilot abstracts many universal interfaces from these requests. First, we have to clarify what an interface is in our paper. Similar to human-defined functions, an interface is a code module consisting of a name, parameters, a functional description, and an implementation code. It performs specific tasks such as data retrieval, computation, and visualization. In our paper, each interface is designed, tested, and optimized iteratively by Data-Copilot.
First, starting from the initial request, we iteratively feed self-exploration requests and related data parsing files into LLMs, prompting the LLM to design complete code modules for request solving. Each code module corresponds to an interface and is stored in the interface library. During each iteration, LLMs are instructed to prioritize utilizing the existing interfaces within the library. If the available interfaces are insufficient for request solving, the LLMs design a new interface.
As shown in Figure 2, for the first request: I want to see the GDP trend over ..., Data-Copilot design two interfaces: Query-GDP(...) and Plot-Line(...). As for the second request: Compare the CPI of China and the USA over ...., since the previous two interfaces could not solve this request, a new interface, Query-CPI(...), is designed by LLMs.
Interface Testing
After designing a new interface for a request (request-i $\rightarrow$ interface-i), Data-Copilot autonomously tests its correctness based on compiler feedback. First, Data-Copilot uses the request-i as the seed to generate massive similar requests as test cases. Then it tests the new interface-i one by one. E.g., in Fig. 2, for newly designed interface-i: Query-GDP(), Data-Copilot mimics request-1 to generate many similar requests to test the Query-GDP(): "I want to see USA’s GDP over the past 10 years", "I want to see China’s GDP for the last year". If the interface passes all test cases, it is retained; otherwise, Data-Copilot utilizes compiler error feedback for self-reflection, correcting the erroneous code snippets until it successfully passes the tests.
Interface Optimization
To further improve the generality of interfaces, Data-Copilot also optimizes similar interfaces. Similar to how human software developers, each time a new interface is designed, this optimization process is triggered: evaluating whether the newly designed interface can be merged with previous ones.
  • Retrieve similar interfaces. After a new interface is designed, we first retrieve the top N similar interfaces from the existing interface library. Specifically, we use gte-Qwen1.5-7B-instruct (Li et al., 2023d) to obtain embeddings of each interface code and then calculate their similarity.
  • Decide whether to merge. Then LLMs are prompted to evaluate the functionality, parameters, and logic of the new interface and the top N similar interfaces, autonomously deciding whether to merge the new interface with existing ones. If the LLM deems no merging necessary, the new interface is retained. As shown in Figure 2, two interfaces Query-GDP(...) and Query-CPI(...) are merged into Query-GDP-CPI(...). This optimization process makes each interface more general and unique, reducing interface redundancy.
  • Test the optimized interface. When two interfaces are merged, Data-Copilot also needs to test the newly merged interface. The test cases from the two original interfaces are used to validate the merged interface. The output of the new interface must be consistent with the two original interfaces in both format and content.
For each request, Data-Copilot alternates the above three steps: interface design, testing, and optimization until all the requests can be covered by these interfaces. As shown in Figure G3, Data-Copilot designs many interfaces for different task and operation types, e.g., data acquisition, prediction, visualization, and DataFrame manipulation. We provide some detailed cases in Section G.1.
2.2 Workflow Deployment
In the first phase, Data-Copilot designs many generic interfaces with different functionalities in advance. They are versatile and error-free, capable of handling the majority of human requests in real-world scenarios. As shown in Figure 2, Data-Copilot only needs to invoke the relevant interfaces for human requests, which is more accurate and efficient (interface workflow). Furthermore, for some long-tail or new requests that cannot be covered by the existing interfaces, Data-Copilot can flexibly generate raw code to handle them (interface-code hybrid strategy) and document these requests for updating the interface library in the future.
Task Selection and Interfaces Retrieval
Considering the significant differences between the interfaces involved in different requests, it is unnecessary to load all interfaces every time. We design a simple yet efficient interface retrieval strategy: hierarchical retrieval. Specifically, we organize all designed interfaces in a hierarchical structure. Each interface is grouped into different tasks (stock task, fund task, etc.) and different operation types (data acquisition, processing, visualization, etc.). Upon receiving a real-time request, Data-Copilot first determines the appropriate task types and required operation types and then loads the interfaces associated with these types for subsequent workflow planning. This can significantly reduce the number of interfaces in the prompt.
Planning Workflow with Different Structures
After reading the descriptions of all retrieved interfaces, Data-Copilot plans an interface workflow to process the user request. A workflow refers to multiple interfaces that are called in a certain order, forming a chain, parallel, or complex loop structures. Therefore, Data-Copilot needs to autonomously determine which interfaces to invoke, in what sequence, and which parameters to populate. The interface invocation workflow is output in JSON format: Interface Call: {step1={"arg":"", "Interface":"", "output":"" ,"desc":""}, step2={..}, ..}. The prompts and cases are shown in Sections F.2 and G.2.
As shown in Figure 2, Data-Copilot designs interface workflows with different structures, e.g., sequential, parallel, or loop. For instance, the first request in Figure 2: "Compare the return of CSI-300, GEM and CSI-1000 this year", Data-Copilot plans a parallel workflow with 5 steps: Data Acquisition(), Calcuate Return() in parallel for calculating three indices, followed by Visualization(). In the second case, it deploys a loop workflow by calling the Loop().
Interface-Code Hybrid Generation
In the real world, it is inevitable to encounter "uncovered" requests that cannot be addressed by existing interfaces. As a remedy, Data-Copilot adopts an interface-code hybrid generation strategy. Specifically, Data-Copilot prioritizes invoking existing interfaces to resolve user requests. It can solve most common user requests. However, if the deployed workflow continues to fail, or if Data-Copilot proactively determines that the current request cannot be resolved by existing interfaces, our system is allowed to generate raw code directly or generate a combination of raw code and interfaces for request solving. This design endows Data-Copilot with the flexibility to handle diverse requests.
Besides, each time Data-Copilot encounters such "uncovered" requests, it also records them in the document. After accumulating a certain number of new requests, we re-initiated the Data Exploration stage (Section 2.1) to develop new interfaces for these newly emerged requests. It empowers Data-Copilot to continuously update its interface library throughout human interactions.
Table 1: Statistics on human-proposed requests and self-exploration dataset.
Source #Cases Form Usage Ratio of 4 Tasks Ratio of 3 Complexities
Human-proposed 173 Query Seed Set 8: 4: 4: 1 2.0: 1.5: 1
Self-Exploration 3000 Query Interface Design 8.7: 3.3: 4.2: 1 1.9: 1.5: 1
Self-Exploration & Human 547 Query, Label Test Set 8.6: 3.3: 4.3: 1 1.8: 1.4: 1
3 Qualitative Evaluation
3.1 Environment and Data Sources
Data-Copilot is evaluating on Chinese financial market data, encompassing massive stocks, funds, economic data, real-time news, and company financial data. Similar to many real-world financial companies, Data-Copilot utilizes data interfaces provided by Tushare [^1] https://tushare.pro/ to access vast amounts of financial data, including time-series data spanning over 20 years for more than 4,000 stocks, funds, and futures. In the first phase, we use a strong LLM (e.g., gpt-4-turbo) for data exploration and a lightweight LLM (e.g., gpt-3.5-turbo) for workflow deployment.
3.2 The Creation of Dataset
Human-proposed Request as Seed. We invite 10 students in economics to submit 50 requests each. These requests cover a wide range of common needs, including stock tasks, fund tasks, and corporate tasks, and also involve different difficulty levels. We treat these 500 human-proposed requests as a distribution of real-world scenarios, where the ratio of four task types is Stock: Fund: Corporation: Other = 8: 4: 4: 1. Then we filter out highly similar requests by computing their sentence embedding. Lastly, we retain 173 high-quality requests, which are used as the seeds for Data Exploration phase of the Data-Copilot.
Self-Exploration Request Set. As mentioned in Section 2.1, we employ self-exploration to expand the request set. Specifically, we feed 173 human-proposed seed requests along with all data descriptions and schema into GPT-4 for data exploration and request synthesis. When self-exploration, we use keywords to control the distribution of the synthesized requests, ensuring that they can closely align with real-world scenarios. In this phase, we generate 6480 requests. After that, we filter out highly similar requests and retain 3547 requests. Then we adopted a stratified sampling strategy to sample 547 instances as the test set from each task type. The remaining 3000 requests are used for interface design. We analyze the number of subsets categorized by four task types (stock, fund, corporation, and others) and three complexity levels (single-entity, multi-entities, and multi-entities with complex relations). The results indicate that our synthesized requests closely align with human distribution. The detailed statistics of three datasets are shown in Tables 1, B1 and B1.
Table 2: Human evaluation on the human-proposed seed set and self-exploration requests.
Number Task Difficulty Request Rationality Expression Ambiguity Answer Accuracy
Seed Set 173 3.5 4.3 4.4 -
Test Set 547 3.9 4.2 4.0 4.8
Annotate Answer Table for Test Set. Human annotators are instructed to annotate data-table as answers for 547 test requests, which is used for benchmarking our system’s output. Specifically, annotators manually retrieve and process the corresponding data based on the testing request, and record the final data table before chart plotting. These data tables are used as answer annotations. For example, for the request: "I want to see China’s GDP trend over the past 5 years", the answer data-table is "[2023: 17.8 trillion, 2022: 17.9 trillion, 2021: 17.8 trillion, 2020: 14.6 trillion, 2019: 14.3 trillion]".
The Quality of Self-exploration Requests. To assess the quality of self-exploration requests, we invited four additional graduate students to manually evaluate our test set (546 requests) and human-proposed requests (173 seed set) on four criteria: task difficulty, request rationality, expression ambiguity, and answer accuracy. We provide detailed guidance for each criterion in Section B.2. The results of two sets are shown in Table 2. We observed that the synthesized requests exhibit a comparable quality to human-proposed requests, with slightly higher difficulty, ambiguity, and similar rationality. It ensures that our test set can reflect most of the real-world demands. Besides, the evaluators gave a high score of 4.8 on the label (answer table) of our test set, which also ensures the accuracy of our dataset.
3.3 Qualitative Results
The whole system is shown in Figure G9, where Data-Copilot dynamically displays the intermediate process (Solving Step) and the final results (bar, chart, text) for the user request. The invocation workflow usually consists of multiple steps, each of which invokes one or multiple interfaces. For example, Data-Copilot addresses the request in Figure 2 ("Compare the return of ...") via a five-step procedure, wherein the first two steps invoke the same interface in parallel.
Data-Copilot designs versatile interfaces through data exploration. After exploring the Chinese financial data and processing 3000 self-exploration requests, Data-Copilot designed a total of 73 interfaces, comprising 5 functionalities and 4 tasks: data acquisition, index calculation, table manipulation, visualization, and general processing logic. These tasks include stocks, funds, corporations, and others. We present some representative interfaces and the statistics of the interfaces in Figure G3. Most interfaces are highly versatile with multiple functions. For example, get-stock-prices-data() interface can retrieve data on a daily, weekly, or monthly basis. These versatile interfaces endow the Data-Copilot with the capability to address complex requests.
Data-Copilot can flexibly deploy interfaces or raw code for diverse demands. As mentioned in  Section 2.2, for some "new" requests, Data-Copilot is capable of addressing them by combining both raw code and existing interfaces. As shown in Figure 2 (3rd example), the user requests a radar chart, which was not directly supported by the available interface. Data-Copilot first leverages the existing interface to retrieve and process the data. Then based on the dataframe produced by the last interface, it generates raw code to render a radar chart. This design grants Data-Copilot greater flexibility to handle the constantly emerging new requests.
Interface Workflow is human-friendly, offering higher interpretability. Upon executing the workflow via the compiler, Data-Copilot produces the visualized results. Also, it provides a comprehensive summary of the entire processing process based on the workflow. As shown in Figures 2, G2, G4, G6, G5 and G7, Data-Copilot’s outputs (workflow, visualized results and summary) are highly intuitive and clear. It encapsulates complex code processing logic into a step-by-step interface invocation workflow, which is easier for humans to understand and inspect.
Table 3: Pass@1 of Data-Copilot and direct code generation (Raw Code) when using different LLMs for workflow deployment: GPT-4o, GPT-4-turbo, GPT-3.5-turbo, Llama-3-70B-chat, Llama-2-70B-chat, Llama-2-13B-chat, Codellama-13B, Vicuna-13B-v1.5).
GPT-4o GPT-4 GPT-3.5 Llama3-70B Llama2-70B Llama2-13B Codellama13B Vicuna
Raw Code 73.6 70.0 28.5 52.2 29.6 17.1 21.0 17.3
Ours 77.9 $\uparrow$4.3 74.6 $\uparrow$4.6 70.2 $\uparrow$41.7 70.9 $\uparrow$18.7 49.3 $\uparrow$19.7 36.9 $\uparrow$19.8 50.8 $\uparrow$29.2 33.2$\uparrow$15
4 Quantitative Evaluation
4.1 Experiments Settings
Baselines and Experiments Details We first compare Data-Copilot with direct code generation (denoted as Direct-Code) using different LLMs. Besides, we also compare Data-Copilot with many prompting and agent strategies, including: 1. ReAct (Yao et al., 2022): LLMs use neural language (thought) and code to resolve the user request step-by-step. 2. Reflexion (Shinn et al., 2023): LLM can reflect on the feedback from the compiler at each step, progressively fulfilling user requests. For the Reflexion strategy, we limit it to a maximum of two rounds of reflection. 3. Multi-agent collaboration (Wu et al., 2023b; Hong et al., 2023; Liang et al., 2023a): We adopt three LLM-based agents, i.e., two coders and a manager, handling user requests simultaneously and outputting the final solution after mutual discussion. For a fair comparison, Data-Copilot employs the same LLM as all baselines for evaluation when invoking pre-designed interfaces. The detailed prompts of all baselines are shown in Section F.3.
Evaluation Methods. All methods are evaluated on the 547 testing instances. We focus on the accuracy of the final results and the efficiency of the response. Considering that the output of most requests is images or charts, we evaluate from two perspectives: whether the data used for plotting is correct and whether the drawn image meets human requirements. We use GPT-4-turbo and GPT-4o to conduct automatic evaluations. (1) Data-table Evaluation: We first obtain the dataframe before visualization, then use GPT-4-turbo to compare this data table with the answer table (labeled by humans in Section 3.2). (2) Image Evaluation: To evaluate whether the final image meets human requests, we feed the human request, human-annotated answer table, and the output image into GPT-4o for evaluation. We instruct GPT-4o to check whether visual elements in the image meet the human requests, including numerical points, lines, axes, image aesthetics, and style. Finally, the results of data and image evaluation are combined as Pass@1. Lastly, we calculate the total number of tokens consumed by each method (denoted as #Token), which represents the response efficiency.
Table 4: We demonstrate the accuracy and efficiency of all methods with gpt-3.5-turbo
Methods Pass@1(%) #Token
Direct-Code 28.5 $\pm$2.1 823
ReAct 44.1 $\pm$3.4 1515
Reflexion 59.0 $\pm$3.9 2463
Multi-Agent 57.4 $\pm$1.8 2835
Data-Copilot 70.2 $\pm$2.5 561.2
Data-Copilot + ReAct 71.5 $\pm$1.7 834
Data-Copilot + Reflexion 71.8 $\pm$2.2 978
Table 5: All samples are divided into 3 complexity levels. We compute Pass@1.
Methods Single Multiple Complex Overall
Direct-Code 42.4% 29.0% 1.6% 28.5%
ReAct 56.3% 48.6% 14.6% 44.1%
Reflexion 67.8% 55.1% 48.1% 59.0%
Multi-agent 63.2% 64.7% 35.5% 57.4%
Data-Copilot 71.8% 70.0% 67.1% 70.2%
4.2 Comparison Results
Data-Copilot Significantly Reduces the Risk of Failure When Real-time Deployment. As demonstrated in Table 3, Compared to directly generating code, Data-Copilot achieves significant improvements across all LLMs, with particularly strong gains of +41.7, +18.7, and +29.2 on GPT-3.5-turbo, Llama3-70B, and Codellama-13B, respectively. By decoupling direct code generation into two stages: interface design in advance and real-time workflow deployment, Data-Copilot can elevate many smaller LLMs (e.g., Lllama3-70b, GPT-3.5) to GPT-4 level during real-time deployment. Through careful comparison, we observe that these pre-designed interfaces significantly reduce the occurrence of errors during real-time responses, as they have already been optimized and validated during the design phase. In contrast, directly generating code from scratch makes it easy to overlook certain details or make mistakes. Besides, we also found that even with a powerful LLM like GPT-4, Data-Copilot also achieves an improvement of over +4%, validating its generalizability.
Data-Copilot Outperforms Advanced Agent Strategies in Both Accuracy and Efficiency. As shown in Table 5, Data-Copilot surpasses all baseline strategies in success rate (Pass@1), and efficiency (#token). Compared to the best baseline (Reflexion), Data-Copilot achieved an +11.2% improvement in accuracy and a -75% reduction in token consumption. Besides, as shown in the last two rows of Table 5, Data-Copilot can seamlessly integrate with agent strategies such as ReAct and Reflexion into the workflow deployment. For example, when combined with ReAct, Data-Copilot invokes an interface, obtains an intermediate result, and then reasons to invoke the next interface, further improving its performance by +1.3.
Data-Copilot Reduces Repetitive Generation in the Real World. In the real world, numerous requests are similar or even repetitive. Data-Copilot can save 70% of the baseline’s consumption because its output is very concise and only contains interface names and arguments. In contrast, baseline strategies have to repetitively generate complete code for each request, consuming many more tokens.
Table 6: Different LLM combinations for two stages. Using a weaker LLM for interface design affects both the interface’s quantity and quality.
Direct Code Generation: GPT-3.5: 28.3, GPT-4: 70
LLM for Workflow Deployment
GPT-3.5 GPT-4
LLM for Interface Design GPT-3.5 31.9 49.6
GPT-4 70.2 74.6
Table 7: We ablated 3 designs from Data-Copilot.
Methods Pass@1(%)
Direct Code Generation 28.5
Data-Copilot 70.2
$w/o$ Self-Exploration Request 35.9
$w/o$ Interface Optimization 42.1
$w/o$ Interface-Code Hybrid 63.8
4.3 Analysis: Why Data-Copilot Can Bring Improvements?
Data-Copilot Exhibits a Pronounced Advantage on More Complex Scenarios. To better analyze the behavior of Data-Copilot, we categorize each test request into different subsets based on the number of entities it involved: single entity, multiple entities, and multiple entities with complex relations (e.g., loop calculations). The statistics for each subset are shown in Section B.1. As shown in Table 5, for the single entity subset (denoted as Single), which represents the simplest case, the baselines and our method perform similarly, with only a 4% improvement. However, for the more challenging multi-entities (denoted as Multiple) or multi-entities with complex relation (denoted as Complex), our improvements are substantial, with enhancements of 5.3% and 19%, respectively. We find that in these complex cases, baseline methods often generate logically incorrect code on the first attempt and may omit some critical steps. In contrast, Data-Copilot only needs to invoke versatile interfaces and deploy workflows, significantly reducing the complexity of real-time responses. These findings highlight the advantages of Data-Copilot, which reduces task complexity by pre-designing interfaces in advance.
Interface Quality is the Key Factor. We analyze the effects of using different LLM combinations for two stages: interface design and workflow deployment. As shown in Table 7, we observed that when using a weaker LLM for interface design, the effectiveness is significantly diminished. For example, using GPT-3.5-turbo (1st stage) and GPT-4-turbo (2nd stage) resulted in a score of only 49.6, which is even lower than directly using GPT-4-turbo for code generation (70.2). This phenomenon can also be seen in other combinations. After manually checking the interfaces in this case, we found these interfaces designed by GPT-3.5-turbo are limited in number and of poor quality. Therefore, we chose to use GPT-4-turbo for interface design and optimization, while employing GPT-3.5-turbo for real-time deployment, achieving a balance between accuracy and efficiency.
4.4 Ablation Study
As shown in Table 7, we ablate Data-Copilot from the following three aspects:
  1. 1.
    Self-exploration Request. We ablate the self-exploration request process Section 2.1. Instead, we directly used seed requests for interface design and optimization. In this case, the performance dropped significantly by 34.3. This may be because there are too few seed requests, which are insufficient to design universe interfaces.
  2. 2.
    Interface Optimization. We ablate the interface optimization phase during interface design. It means every successfully designed interface was retained in the interface library. We observed that performance was also significantly affected (-28 points). Without interface optimization, there are many similar interfaces, which could hurt the effect of workflow invocation.
  3. 3.
    Interface-Code Hybrid Generation. Data-Copilot only can invoke interfaces during workflow deployment. Directly generating code is not allowed even when encountering requests that cannot be resolved by interfaces. After ablation, we observe a performance drop of 6.4 points. These losses are mainly caused by new requests that cannot be covered by existing interfaces. It highlights the flexibility of Data-Copilot, which can flexibly solve familiar requests by invoking interfaces or handle new requests by generating code directly.
5 Conclusion
We propose a data analysis agent, Data-Copilot, to address a wide array of data-related tasks. Data-Copilot is a code-centric agent, which generates codes to process large-scale data, and generates generic interface modules in advance through a data exploration process to improve the success rate of real-time requests. It autonomously designs universal interface modules for various types of data and potential user demands and invokes these interfaces in real-time response for more reliable problem-solving. Our experiments show that Data-Copilot flexibly handles challenging requests about Chinese financial data, with a higher success rate and less token consumption.
Appendix
Appendix A Preprocessing of Workflow Deployment
Intent Analysis
To accurately comprehend user requests, Data-Copilot first parses the time, location, data object, and output format of user requests, which are critical to data-related tasks. For example, the request is: "I want to compare the GDP and CPI trend in our area over the past five years", Data-Copilot parses it as: "Draw a line chart of China’s national GDP and CPI per quarter from May 2019 to May 2024 for comparison". To achieve this, we first invoke an external API to obtain the local time and network IP address, then feed this supportive information into LLMs along with the original request to generate the parsed result.
Multi-form Output
Upon execution of the workflow, Data-Copilot yields the desired results in the form of graphics, tables, and descriptive text. Additionally, it also provides a comprehensive summary of the entire workflow. It greatly enhances the interpretability of the whole process, as the interface workflow is easy for humans to read and inspect. As the example shown in Figure G2, the request is "Forecasting China’s GDP growth rate...". Data-Copilot first interprets the user’s intent. Then it deploys a three-step workflow: 1) Invoking get-GDP-data() interface to acquire historical GDP data. 2) Invoking predict-next-value() interface for forecasting. 3) Visualizing the output.
Appendix B Evaluation Details
B.1 The Definitions of Three Complexity Levels for Testset
We categorize our test set into three types based on the number of entities involved: single entity, multiple entities, and multiple entities with complex relations (e.g., loop calculations). The ststics of three subset are shown in Table B1.
  • Single entity: Requests involve a single entity and can be resolved step-by-step. E.g., "query based on a specific condition".
  • Multiple entities: Requests require processing multiple entities simultaneously. E.g., "compare a certain metric across multiple entities".
  • Multiple entities with complex relations: Involving multiple entities and containing loops, nesting, and other intricate logic. E.g., "List the top 10 stocks by yesterday’s price increase that also in the internet industry."
Table B1: Statistics of four task types and three complexity levels on our test set.
Task Types Request Complexity
Single Entity Multiple Entities Multi-entities with Complex Relation Overall
Stock 79 106 87 272
Fund 55 36 22 113
Corp. 97 30 4 131
Other 6 12 13 31
Total 237 184 126 547
B.2 Human Evaluation on Test Benchmark
We invite four more graduate students to manually evaluate our benchmark according to four criteria: task difficulty, request rationality, expression ambiguity, and answer accuracy. The belief guidance for human evaluation is as follows:
  • Task Difficulty: The difficulty of the task, whether it requires multiple steps… Scoring Criteria: 5: very difficult,…, 1: easy.
  • Request Rationality: Whether the request is reasonable, or if it is strange and does not align with human habits… Scoring Criteria: 5: Reasonable, aligns with humans..
  • Expression Ambiguity: Whether the phrasing of the request is ambiguous, or if the entities involved are unclear… Scoring Criteria: 5: Clear without any ambiguity…
  • Answer Accuracy: Whether the answer table is correct, detailed, and comprehensive, totally meeting user expectations… Scoring Criteria: 5: Data is totally correct … 3: Partial…
Figure B1: We count the keywords for each request type (Stock, Corp., Fund, Others) in the test set.
B.3 Expanding to Other Programming Languages.
In addition to the Python language, Data-Copilot exhibits excellent scalability, allowing for easy switching to other programming languages by simply regenerating the corresponding interfaces. We tested three programming languages: Python, C++, and Matlab. The results indicate that Python performed the best (Pass@1: 70.2), followed by C++ (54.2), with Matlab (36.5) yielding the poorest results. Upon examination, we found that Matlab code often suffers from formatting errors or dimensional discrepancies in the data, rendering the program non-executable. We speculate this may be related to a lack of sufficient Matlab code in the pre-training corpus. So ultimately, we opted for Python for Data-Copilot.
Appendix C Limitations
Data-Copilot proposes a new paradigm for addressing the data-related task, through LLM. But we want to highlight that it still remains some limitations or improvement spaces:
1) Online Design Interface. The essence of Data-Copilot lies in effective interface design, a process that directly affects the effectiveness of subsequent interface deployments. Currently, this interface design process is conducted offline. Therefore, it is crucial to explore how to design the interface online and deploy it simultaneously. It greatly broadens the application scenarios of Data-Copilot.
2) System stability The interface deployment process can occasionally be unstable. The main source of this instability is because LLM is not fully controllable. Despite their proficiency in generating the text, LLMs occasionally fail to follow the instructions or provide incorrect answers, thus causing anomalies in the interface workflow. Consequently, finding methods to minimize these uncertainties during the interface dispatch process should be a key consideration in the future.
3) Data-Copilot possesses the potential to handle data from other domains effectively. Currently, our focus is on developing an LLM-based agent for the data domain. Due to the limited access to data, we chose the China Financial Data, which includes stocks, futures, finance, macroeconomics, and financial news. Although these data all belong to the financial domain, the data volume is extremely large, and the data schemas are highly different. The corresponding user’s requests are also diverse, which poses a great challenge to the current LLM. However, Data-Copilot has adeptly accomplished this task. Therefore, we believe Data-Copilot also possesses the potential to effectively handle data from other domains.
Appendix D Related Works
In the recent past, breakthroughs in large language models (LLMs) such as GPT-3, GPT-4, PaLM, and LLaMa [Brown et al., 2020, Chowdhery et al., 2022, Zhang et al., 2022, Zeng et al., 2023, Touvron et al., 2023, Ouyang et al., 2022, OpenAI, 2023, Wei et al., 2022a] have revolutionized the field of natural language processing (NLP). These models have showcased remarkable competencies in handling zero-shot and few-shot tasks along with complex tasks like mathematical and commonsense reasoning. The impressive capabilities of these LLMs can be attributed to their extensive training corpus, intensive computation, and alignment mechanism [Ouyang et al., 2022, Wang et al., 2022b, a].
LLM-based Agent Recent studies have begun to explore the synergy between external tools and large language models (LLMs). Tool-enhanced studies  [Schick et al., 2023, Gao et al., 2022, Qin et al., 2023a, Hao et al., 2023, Qin et al., 2023b] integrate external tools into LLM, thus augmenting the capability of LLMs to employ external tools. Several researchers have extended the scope of LLMs to include the other modality [Wu et al., 2023a, Surís et al., 2023, Shen et al., 2023, Liang et al., 2023b, Huang et al., 2023]. In addition, there are many LLM-based agent applications [Xie et al., 2023], such as CAMEL [Li et al., 2023a], AutoGPT [^2] https://github.com/Significant-Gravitas/Auto-GPT, AgentGPT [^3] https://github.com/reworkd/AgentGPT, BabyAGI [^4] https://github.com/yoheinakajima/babyagi, BMTools [^5] https://github.com/OpenBMB/BMTools, LangChain [^6] https://github.com/hwchase17/langchain, Agentverse [Chen et al., 2023b], Autoagent [Chen et al., 2023a], MetaGPT [Hong et al., 2023], AutoGEN [Wu et al., 2023b], etc. Most of them are focused on daily tools or code generation and do not consider the specificity of data-related tasks. Except for learning to operate the tools, several contemporaneous studies [Cai et al., 2023, Qian et al., 2023] have proposed to empower LLMs to create new tools for specific scenarios like mathematical solving and reasoning. These impressive studies have revealed the great potential of LLM to handle specialized domain tasks.
Applying LLM To Data Science Apart from these studies, the application of large models in the field of data science has garnered significant interest among researchers. FLAME [Joshi et al., 2023] investigates the feasibility of using NLP methods to manipulate Excel sheets. StructGPT [Jiang et al., 2023] explore reasoning abilities of LLM over structured data. LiDA [Dibia, 2023] and GPT4-Analyst [Cheng et al., 2023] focus on automated data exploration. Besides, many reseraches [Liu et al., 2023, Chang and Fosler-Lussier, 2023, Dong et al., 2023], like Sheet-Copilot [Li et al., 2023b], BIRD [Li et al., 2023c], DAIL-SQL [Gao et al., 2023], DIN-SQL [Pourreza and Rafiei, 2023], PET-SQL [Li et al., 2024b], DB-Copilot [Wang et al., 2023b], MAC-SQL [Wang et al., 2023a], ACT-SQL [Zhang et al., 2023], ChatBI [Lian et al., 2024], CodeS [Li et al., 2024a], SQLPrompt [Sun et al., 2023a], ChatDB [Hu et al., 2023], SQL-PaLM [Sun et al., 2023b], and DB-GPT [Xue et al., 2023], EHRAgent [Shi et al., 2024] apply LLMs to Text2SQL and table rasoning. Chain-of-Table [Wang et al., 2024] proposes a step-by-step reasoning strategy based on the table.
Distinct from these approaches, our Data-Copilot system is different in the following aspects:
(1) Different Task Settings: Unlike Text2SQL methods, Data-Copilot focuses more on how to utilize data and present it according to user preferences, rather than retrieving data from databases. We assume that Data-Copilot has access to a vast amount of data, which is available via API.
(2) Richer Usage Scenarios: Data-Copilot needs to meet a diverse range of human needs, such as forecasting and plotting. Therefore, it employs programmable language like Python for data processing, offering more varied use cases than SQL.
(3) Distinct Implementation Approach: Data-Copilot is a code-centric agent for data processing, rather than relying on SQL or pre-existing tools. This approach is more flexible and can meet complex user requirements.
Appendix E Visualization
We provide several cases in this section to visualize workflow deployed by Data-Copilot, which includes queries about diverse sources (stocks, company finance, funds, etc.) using different structures (parallel, serial, and loop Structure).
Different Structures
As shown in Figure G4,G5, Data-Copilot deploys different structural workflows based on user requirements. In Figure G4, the user proposes a complex request, and Data-Copilot deploys a loop structure to implement the financial data query of each stock, and finally outputs a graph and table in parallel. In Figure G5, Data-Copilot proposes a parallel structure workflow to meet user request (the demand for comparison in user request) and finally draws two stock indicators on the same canvas. These concise workflows can cope with such complex requests well, which suggests that the data exploration and workflow deployment process of Data-Copilot are rational and effective.
Diverse Sources
Figure G6, G7, G8 demonstrate that Data-Copilot is capable of handling a large number of data sources, including stocks, funds, news, financial data, etc. Although the formats and access methods of these data types are quite different, our system efficiently manages and displays the data through its self-designed versatile interface, requiring minimal human intervention.
Appendix F Detailed Prompts
We provide detailed prompts for our Data-Copilot and baselines. Specifically, for the data exploration, we provide detailed prompts in Section F.1 for four phases: self-exploration, interface design, and interface optimization. For the workflow deployment phase, we also provide prompts for three key procedures in Section F.2: Intent Analysis, Task Selection, and Planning Workflow. Additionally, we outline prompts for all baselines in Section F.3: Direct-Code, ReAct, Reflection, and Multi-Agent Collaboration strategies.
F.1 Prompts For Data Exploration
Explore Data by self-exploration Phase
###Instruction: Given some data and its description, please mimic these seed requests and generate more requests. The requests you generate should be as diverse as possible, covering more data types and common needs.
###Seed Request: {request1, request2,...}
###parsing file for GDP-Data:{
Description: This data records China’s annual and quarterly GDP...,
Access Method: You can access the data by pro.cn-gdp(start-time, end-time, frequency,...), start-time means....,
Output Schema: The data return 9 columns, including quarter: quarter, gdp: cumulative GDP, gdp-yoy: quarterly Year-on-Year growth rate, pi,...,
Usage:{
Example: pro.cn-gdp(start-q=’2018Q1’, end-q=’2019Q3’, frequency=’quarter’),
First Row: {2019Q4 990865.1 6.10, ...},
Last Rows: {2018Q4 900309.5 6.60, ...}}}
###parsing file for stock_data: {...}
Interface Design:
###Instruction: You are an experienced program coder. Given a request and some existing interfaces, you should use these interfaces to solve the request or design new interfaces to resolve my request.
(1) You should define the name, function, inputs, and outputs of the interface. Please describe the functionality of the interface as accurately as possible and write complete implementation code in the new interface.
(2) Finally please explain how to resolve my request using your newly designed interfaces or existing interfaces in the neural language.
###Output Format:
Your newly designed interfaces:
Interface1={Interface Name: {name},
Function description: {This interface is to ...},
Input: {argument1: type, argument2: type, ...},
Output: {pd.DataFrame} }
Interface2=....
The solving process for request: {To fullfil this request, I design a interface ...}
###The user request: {input request}
###Data Files: {All data parsing files}
###The existing interfaces: {all interfaces in library}
Interface Optimization
###Instruction: Please check that the interface you have designed can be merged with any existing interfaces in the library.
(1) You should merge interfaces with similar functionality and similar input and output formats into a new interface.
(2) You can use parameters to control the different inputs, if you want to merge two interfaces.
(3) Please explain your reason for merging and output all interfaces in the library after merging.
(4) If you dont think a merge is necessary, then just add new interfaces into the existing interface library and output them all.
###New interfaces: {interfaces}
###Existing interfaces: {interface1, interface2, interface3, ..}
###Output Format:
The reasons for merging: {reason}
Interfaces after merging: {interface1, interface2, optimized interface3, ..}
F.2 Prompts For Workflow Deployment
Intent Analysis Phase
### Analysis Prompt: Please parse the input request for time, place, object, and output format. You should rewrite the instruction according to todays date. The rewritten new instruction must be semantically consistent and contain a specific time and specific indicators.
### Output Format: Rewritten Request. (Time:%s, Location:%s, Object:%s, Format:%s).
### User Request: Today is {Timestamp}. The user request is {Input Request}.
Please output a Rewritten Request.
Task Selection
###Select Prompt: Please select the most suitable task according to the given Request and generate its task_instruction in the format of task={task_name: task_instruction}. There are four types of optional tasks. [fund_task]: used to extract and process tasks about all public funds. [stock_task]: for extracting and processing tasks about all stock prices, index information, company financials, etc., [economic_task]: for extracting and processing tasks about all Chinese macroeconomic and monetary policies, as well as querying companies and northbound funds, [visualization_task]: for drawing one or more K-line charts, trend charts, or outputting statistical results.
###Output Format: task1={%s: %s}, task2={%s: %s}
###User Request: {Rewritten Request}.
Please output a task plan for this request.
Planning Workflow
###Planning prompt: Please use the given interface (function) to complete the Instruction step by step. At each step you can only choose one or more interfaces from the following interface library without dependencies, and generate the corresponding arguments for the interface, the arguments format should be strictly in accordance with the interface description. The interface in the later steps can use results generated by previous interfaces.
###Output Format:
Please generate as json format for each step:step1={"arg1": [arg1,arg2...], "function1": "%s", "output1": "%s", "description1": "%s"}, step2={"arg1": [arg1,arg2..], "function1": "%s", "output1": "%s", "description1": "%s"}, ending with ###.
###User Request: {Task Instruction}.
Please output an interface invocation for this instruction.
F.3 Prompts For Baselines
Direct-Code LLM
###Instruction: You are an artificial intelligence assistant. Given some data access methods and a user request, you should write a complete Python code to fulfill the users request. Your code must completely fulfill all the users requirements without syntax errors!
###User Request: {User request}
###Data files: {All data files}
Please solve the request by Python Code.
Step-by-Step ReAct
###Instruction: You are an artificial intelligence assistant. Given some data access methods and a user request, please think step by step and generate your thoughts and actions for each step, and then finally realize the users request.
###User Request: {User request}
###Data files: {All data files}
###Thought Prompt: Please think about the next action that should be taken to handle the user request.
### {Thought: I need to ...}
###Action Prompt: Based on your previous thoughts, please generate a complete Python code to accomplish what you just planned.
### {Action: def get-data()....}
###Observation Prompt: Please summarize the results of the code execution just now and think about whether this result accomplishes what you planned for this step.
### {Action: Yes, I observed that this function successfully fetched the data...}
Step-by-Step Reflexion
###Instruction: You are an artificial intelligence assistant. Given some data access methods and a user request, please think step by step and then generate your thoughts and actions for each step. After the execution of your current action, you need to reflect on the results until your current plan has been successfully completed. Then you think about the next step and then generate your next action, and finally realize the users request.
###User Request: {User request}
###Data files: {All data files}
###Thought Prompt: Please think about the next action that should be taken to handle the user request.
### {Thought: I need to ...}
###Action Prompt: Based on your previous thoughts, please generate a complete Python code to accomplish what you just planned.
### {Action: def get-data()....}
###Observation Prompt: Please record the compilers return results just now and think about whether this result accomplishes what you planned for this step.
### {Action: No, I observe that the compiler returns an error...}
### Reflection Prompt: Please reflect on the error returned by the compiler and regenerate a new Python code to resolve the issue. If the compilation passes without any errors, reflect on whether the current result is what you planned to do.
### {Action: I revise my solution as follows: def get-data2()....}
Multi-agent collaboration
###Instruction For Manager: You are the manager of the project team and you need to lead your team to fulfill user requests. You have two experienced programmers under you (Programmer-A and -B) and you need to assign them the same or different tasks according to the user request, then organize the discussion, and finally solve the problem.
###Instruction For Agent1/Agent2: You are an experienced programmer. Your team has a colleague who is also a programmer and a manager. You need to write code according to the managers arrangement, discuss it with them, improve your program, and get a consensus conclusion.
###User Request: {User request}
###Data files: {All data files}
-------- Phase1 Discussion for Task Assignment -------
###Task Assignment Prompt For 3 Agents: Now lets start discussing how to solve user problems. The speaking order is manager, A and B. After a round of discussion, the manager decides on a preliminary solution and assigns specific tasks to the two programmers
###Manager Response: {To fulfill the request...}
###Programmer A Response: {Yes, I agree with the manager. We should first,....}
###Programmer B Response: {That’s a good solution. We also should...}
###Manager Response: {After discussion, Programmer A should..., B should ....}
-------- Phase2 Coding and Discussion-------
###Prompts For 3 Agents: Two programmers complete their respective tasks as assigned by their manager. Then the manager organizes a discussion on whether the two options and their running results are reasonable. Lastly, two programmers need to regenerate the code.
###Programmer A Response: {Solution1 Generated by A}
###Programmer B Response: {Solution1 Generated by B}
###Prompts For 3 Agents: Please evaluate two solutions. {Solution1, Result1, Solution2, Result2}
###Programmer A Evaluation: {I think ...}
###Programmer B Evaluation: {I agree with ...}
###Programmer A Regeneration: {Revised Solution1}
###Programmer B Regeneration: {Revised Solution2}
-------- Phase3 Discussion-------
-------- Phase4 Final Decision-------
###Prompts For manager: Please output a complete and correct solving code for user request based on the previous discussions of the group.
###Manager Decision: {Final Solution}
Appendix G Case Study
Detailed cases of data exploration and workflow deployment are presented in Sections G.1 and G.2.
G.1 Cases for Data Exploration
---- The First Request: Interface Design -----
Input Request: Get the daily opening price data of Bank of Hangzhou from Jan 23, 2018 to Mar 13, 2019 and plot the trend.
Response: First, we need to design a function to get the opening price data from the banks database. Then, we need another function to process this data and prepare them for graphical representation. Finally, we need a function to plot the trend chart.
New Design Interfaces:
{Interface1=Name:getBankData(), Function:.., Input:.., Output:..}
{Interface2=Name:prepareDataForPlotting(), Function:.., Input:.., Output:..}
{Interface3=Name:plotData(), Function:.., Input:.., Output:..}
The solving process for request: These three new interface functions will meet your needs. In your program, you would first get the data using getBankData, then process the data using prepareDataForPlotting, and finally use plotData to plot the trend.
Complete code: {Complete code using three new interfaces...}
---- The Second Request: Interface Definition Phase-------
Input Request: Get the monthly closing price data of Ningde Times from January 23, 2018 to March 13, 2019, and plot the trend.
Response: To fulfill this requirement, we design a new interface function. In addition, I need to use two existing interfaces to fulfill your requirement.
New Designed Interfaces:
{Interface4=Name: getStockData()..}
Existing Interfaces
{Interface1=Name:getBankData(), ...}
{Interface2=Name:prepareDataForPlotting(), ...}
{Interface3=Name:plotData(), ...}
The solving process for request: You need to call these functions to meet your needs. First, use new interface {getStockData()} to get the data of Ningde Times, then use existing interface {prepareDataForPlotting()} to process this data, and finally use {plotData()} to plot the trend...
Complete code: {Complete code using three new interfaces...}
---- Interface Optimization Phase -------
The reason for merging: Yes, these interface functions can be merged and reformulated to eliminate duplicate functions. We can merge getBankData() and getStockData() into a more generic function get-financial-data(), which can fetch any financial data, including bank and stock data. In addition, we can keep the prepareDataForPlotting and plotData functions, as they can also be applied to different types of financial data]
{Interface1=Interface Name:get-financial-data(), ...}
{Interface2=Interface Name:prepareDataForPlotting(), ...}
{Interface3=Interface Name:plotData(), ...}
G.2 Cases for Real-time Request
### Request: Today is 2019-03-13. The user request is to please help me plot the stock price trend of Guizhou Maotai from January 23, 2018 to today.
### Rewritten Request: Please help me plot the closing price of Guizhou Maotai from Jan 23, 2018 to Mar 13, 2019 (Time:20180123-20190313, Location: China, Object: Guizhou Maotai stock price, Format: Line).
###Plan: task1={"stock_task": "Get the time series data of Guizhou Maotais stock price from 20180123 to 20190313"}, task2={"visualization_task": "Plot the stock price of Guizhou Maotai from 20180123 to 20190313 on a line chart"}
###Function Call:
step1={
"arg1": ["Guizhou Maotai", "20180123", "20190313", "daily"],
"function1": "get_stock_prices_data",
"output1": "result1",
"description1": "Guizhou Maotai daily stock price (open, close, high, low, etc.) time series data"},
step2={
"arg1": ["result1","close"],
"function1": "calculate_stock_index",
"output1": "result2v,
"description1": "Guizhou Maotais closing price time series data"},
step3={
"arg": ["result2", null, "line", "Guizhou Maotai closing price chart from January 23, 2018 to March 13, 2019"], "function": "plot_stock_data",
"output": "result1",
"description": "Guizhou Maotai Stock Price Line Chart"
Figure G2: Data-Copilot deploys workflows to solve users’ prediction request. It invokes three interfaces step by step and generates arguments for each interface.
Figure G3: We have listed some of Data-Copilot’s self-design interfaces, including five categories. Most of the interfaces are a combination of several simple interfaces. For example, the text marked in blue indicates that the interface includes multiple functions.
Figure G4: For complex requests about stock financial data, Data-Copilot deploys a loop workflow to solve user requests and finally outputs images and tables in parallel.
Figure G5: Data-Copilot acquires the P/E data of two stocks in parallel and plots them in a graph.
Figure G6: Data-Copilot also has access to fund data and can query the returns of all funds managed by the fund manager.
Figure G7: Data-Copilot can plot multiple indicators in a single graph by deploying workflows in parallel.
Figure G8: Data-Copilot can provide the latest financial news and company information by deploying the corresponding workflows.
Figure G9: The user interface of our system. The green box (A) is the user input panel, and the purple (B) and red parts (C) are the results returned by the system.